# SCAM: A Real-World Typographic Robustness Evaluation for Multimodal Foundation Models

### v6

@@ 6 Discussion
Our evaluations on the Subtle Character Attacks on Multimodal Models (SCAM) datasets confirm previous findings [10, 11, 12] that multimodal foundation models remain vulnerable to typographic attacks, whether real-world (SCAM) or synthetic (SynthSCAM). Notably, we find that synthetic attacks are just as effective as real-world ones, leading to comparable performance drops.
@@
Despite incorporating a different, “safer” prompt in section˜A.5, a central limitation of our evaluation remains the reliance on “naive prompting” [55, 56] for assessing LVLMs. Adopting more advanced multi-prompt evaluation strategies may yield a more reliable measure of model robustness against typographic attacks. A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility.
@@
As multimodal models are increasingly deployed to safety-critical applications [15, 14, 16, 16], ensuring robustness against typographic attacks is essential to prevent serious consequences. Approaches such as alignment training [57] or (mechanistic) defense techniques [44, 58] offer promising directions for enhancing robustness. Our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder. However, our initial experiments (section˜A.6) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively. Future work will extend this by systematically retraining LLaVA models with a range of vision backbones to evaluate their role in attack susceptibility and to assess whether encoder strength alone suffices or if additional adaptation steps are necessary.
@@ A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
A.2 Zero-Shot Evaluation of LVMs: Text Templates
To compute the cosine similarity between an image and text embeddings in our zero-shot evaluation of Vision-Language Models (VLMs), we generate text embeddings for both the object label and the attack word using multiple textual variations. Specifically, we apply 32 diverse text templates as in [1], each incorporating the object label resp. attack word in a different context. The final embedding is obtained by averaging the embeddings of all templates. The templates are as follows: ( ’a bad photo of a {}.’, ’a photo of many {}.’, ’a sculpture of a {}.’, ’a photo of the hard to see {}.’, ’a low resolution photo of the {}.’, ’a rendering of a {}.’, ’graffiti of a {}.’, ’a bad photo of the {}.’, ’a cropped photo of the {}.’, ’a tattoo of a {}.’, ’the embroidered {}.’, ’a photo of a hard to see {}.’, ’a bright photo of a {}.’, ’a photo of a clean {}.’, ’a photo of a dirty {}.’, ’a dark photo of the {}.’, ’a drawing of a {}.’, ’a photo of my {}.’, ’the plastic {}.’, ’a photo of the cool {}.’, ’a close-up photo of a {}.’, ’a black and white photo of the {}.’, ’a painting of the {}.’, ’a painting of a {}.’, ’a pixelated photo of the {}.’, ’a sculpture of the {}.’, ’a bright photo of the {}.’, ’a cropped photo of a {}.’, ’a plastic {}.’, ’a photo of the dirty {}.’, ’a jpeg corrupted photo of a {}.’, ’a blurry photo of the {}.’ ).
A.3 Model Accuracies on Typographic Attack Datasets
Table 6: Performance of VLMs and LVLMs available through OpenCLIP [20] resp. ollama [21], and Anthropic’s resp. OpenAI’s API on SCAM, PAINT [42], and RTA-100 [44] datasets.
@@ A.4 Architecture Comparison
Table 7: Models chosen from the OpenCLIP suite for architecture comparison in table˜5(a). The table groups models by architecture category and indicates the corresponding fixed training data. In OpenCLIP, v1 refers to LAION-COCO NLLB [51].
@@
A.5 Prompt Variation
Our evaluation of LVLMs in section˜4 focuses on a single prompt. Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation [59, 50], we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt. Contrary to the expectation formulated in these works, this modification does not reduce vulnerability. As illustrated in fig.˜8, the susceptibility to typographic attacks remains largely unchanged.
@@ A.5 Prompt Variation
A.6 Replacing VLM in LLaVA
Previous work [40, 60] suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact. Further, [61] demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in [60].
Our evaluations on SCAM support the conclusion that some VLMs exhibit higher typographic robustness than others. In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA [14, 30, 31] – show heightened vulnerability. To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder [62, 63], which was trained on datacomp1b and differs primarily in training data and regime, not architecture. Compared to the OpenAI encoder, the CLIPA model exhibits a substantially lower robustness drop: only 25% from NoSCAM to SCAM, versus 65% (see table˜1). It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP [48] [^3] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
@@ A.6 Replacing VLM in LLaVA
Table˜8 shows that our CLIPA-based LLaVA variant does not outperform the original model on SCAM, despite using a VLM that is more robust in isolation. Performance on other benchmarks evaluated via VLMEvalKit [65] also drops. These findings suggest that the benefits of a more robust vision encoder do not transfer under naive LLaVA integration.
@@
Table 8: Performance comparison of the original LLaVA-v1.5-7B and our CLIPA-based LLaVA variant across diverse benchmarks. Despite the CLIPA vision encoder showing stronger robustness in isolation, this does not translate to improved performance on SCAM. Evaluations for the upper section are conducted using VLMEvalKit [65]. Other results follow the protocol outlined in the main part of the paper.
Benchmark LLaVA-v1.5-7B (original) LLaVA-v1.5-7B-CLIPA (ours)
MMBench v1.1 [66] 64.32 60.22
MME Perception [67] 1444.81 1378.96
SEED-Bench-1 (img) [68] 66.81 64.02
TextVQA [69] 46.93 45.57
OCRBench [70] 31.90 31.60
OCR-VQA [71] 60.64 57.39
@@
PAINT [42] 52.73 53.64
RTA100 [44] 48.30 47.30
A.7 Typographic Robustness vs. Typographic Understanding
The main paper shows that increasing the size of LLM backbones reduces susceptibility to typographic attacks (see results for LLaVA and Gemma3 in table˜1). Simultaneously, benchmarks such as TextVQA and OCRBench, reported on the VLMLeaderboards [^5] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, indicate that typographic understanding – essentially, reading – also improves with model scale (see table˜9). Together, these findings suggest that typographic robustness emerges as model size increases, alongside improved textual perception. However, [36] reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
Table 9: Performance of selected LLaVA-Next [31] models on three OCR-related benchmarks (TextVQA [69], OCRBench [70] (from VLMEvalKit [65]), and OCR-VQA [71]) shows that increasing LLM backbone size leads to improved typographic understanding. All models use the ViT-L-14-336 vision encoder. Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit [65].

### v7

@@ 6 Discussion
Our evaluations on the Subtle Character Attacks on Multimodal Models (SCAM) datasets confirm previous findings (Goh et al., 2021; Cheng et al., 2024c; Qraitem et al., 2024) that multimodal foundation models remain vulnerable to typographic attacks, whether real-world (SCAM) or synthetic (SynthSCAM). Notably, we find that synthetic attacks are just as effective as real-world ones, leading to comparable performance drops.
@@
Despite incorporating a different, “safer” prompt in section˜A.6, a central limitation of our evaluation remains the reliance on “naive prompting” (He et al., 2024; Cao et al., 2024a) for assessing LVLMs. More advanced multi-prompt evaluation strategies may yield a more reliable measure of robustness against typographic attacks. A further limitation is our exclusive focus on English-language attacks; future work should investigate whether language-specific factors influence model susceptibility..
@@
As multimodal models are increasingly deployed to safety-critical applications (Li et al., 2023b; Hartsock and Rasool, 2024; Zhou et al., 2024), ensuring robustness against typographic attacks is essential to prevent serious consequences. Approaches such as alignment training (Afzali et al., 2024) or (mechanistic) defense techniques (Azuma and Matsui, 2023; Hufe et al., 2025) offer promising directions for enhancing robustness. The aligned NoSCAM/SCAM pairs are well-suited for causal and mechanistic analyses, as they allow for controlled interventions while keeping scene semantics fixed. We view SCAM primarily as a dataset and benchmark to support such work and expect future research to leverage this alignment to develop and evaluate mechanistic defenses against typographic attacks. Furthermore, our evaluations suggest that the vulnerability of LVLMs to typographic attacks is tied to their vision encoder. However, our initial experiments (section˜A.7) show that simply swapping LLaVA’s vision encoder for a more robust one does not reliably transfer robustness when integrated naively. Future work will extend this by systematically retraining LLaVA models with a range of vision backbones to evaluate their role in attack susceptibility and to assess whether encoder strength alone suffices or if additional adaptation steps are necessary.
@@ 7 Conclusion
Broader Impact Statement
Our work highlights a concrete vulnerability in (Large) Vision-Language Models: typographic attacks that cause models to overly rely on embedded text rather than visual content. This vulnerability has clear negative implications. Malicious actors could exploit our findings to systematically uncover weaknesses in deployed vision–language systems, using typographic cues to mislead models in safety-critical settings such as medical decision support, navigation, or content moderation. By releasing a large and diverse dataset of such attacks, we also create the possibility that others could study these loopholes with the explicit aim of designing more effective attacks.
At the same time, we believe the benefits for research and defense outweigh these risks. Robustness research requires shared realistic benchmarks and reproducible baselines. SCAM provides a standardized resource that allows the community to evaluate typographic vulnerabilities and compare defense methods systematically, and to study how training data and model design affect susceptibility. By demonstrating that synthetic attacks reliably mirror real-world ones, we also reduce the need for ongoing collection of handwritten adversarial examples, lowering barriers for future research.
@@ A.1 Example Images of SCAM, NoSCAM, and SynthSCAM
A.2 Example Images of PAINT and RTA-100
Although qualitative inspection is necessarily subjective and 20 samples from the PAINT (Ilharco et al., 2022) resp. RTA-100 (Azuma and Matsui, 2023) datasets in figs.˜9 and 9 cannot capture their full variability, the quantitative evidence in LABEL:tab:dataset_comparison and the attack-word distribution shown in fig.˜3 clarify the structural differences. PAINT and RTA-100 cover comparatively restricted sets of objects, attack words, and scene conditions. SCAM, by contrast, comprises 1162 images with 660 objects and 206 attack words collected by nine contributors across diverse indoor and outdoor environments. The resulting variation in lighting, backgrounds, camera quality, handwriting styles, and viewpoints is visible in fig.˜2, even if such diversity is difficult to discern from limited qualitative examples alone.
Figure 8: 20 random example images from the PAINT (Ilharco et al., 2022) dataset. Best viewed zoomed in.
Figure 9: 20 random example images from the RTA-100 (Azuma and Matsui, 2023) dataset. Best viewed zoomed in.
A.3 Zero-Shot Evaluation of LVMs: Text Templates
To compute the cosine similarity between an image and text embeddings in our zero-shot evaluation of Vision-Language Models (VLMs), we generate text embeddings for both the object label and the attack word using multiple textual variations. Specifically, we apply 32 diverse text templates as in (Radford et al., 2021), each incorporating the object label resp. attack word in a different context. The final embedding is obtained by averaging the embeddings of all templates. The templates are as follows: ( ’a bad photo of a {}.’, ’a photo of many {}.’, ’a sculpture of a {}.’, ’a photo of the hard to see {}.’, ’a low resolution photo of the {}.’, ’a rendering of a {}.’, ’graffiti of a {}.’, ’a bad photo of the {}.’, ’a cropped photo of the {}.’, ’a tattoo of a {}.’, ’the embroidered {}.’, ’a photo of a hard to see {}.’, ’a bright photo of a {}.’, ’a photo of a clean {}.’, ’a photo of a dirty {}.’, ’a dark photo of the {}.’, ’a drawing of a {}.’, ’a photo of my {}.’, ’the plastic {}.’, ’a photo of the cool {}.’, ’a close-up photo of a {}.’, ’a black and white photo of the {}.’, ’a painting of the {}.’, ’a painting of a {}.’, ’a pixelated photo of the {}.’, ’a sculpture of the {}.’, ’a bright photo of the {}.’, ’a cropped photo of a {}.’, ’a plastic {}.’, ’a photo of the dirty {}.’, ’a jpeg corrupted photo of a {}.’, ’a blurry photo of the {}.’ ).
A.4 Model Accuracies on Typographic Attack Datasets
Table 6: Performance of (Large) Vision-Language Models available through OpenCLIP (OpenCLIP, ) resp. ollama (Ollama, ), and Anthropic’s resp. OpenAI’s API on SCAM, PAINT (Ilharco et al., 2022), and RTA-100 (Azuma and Matsui, 2023) datasets.
@@ A.5 Architecture Comparison
Table 7: Models chosen from the OpenCLIP suite for architecture comparison in table˜5(a). The table groups models by architecture category and indicates the corresponding fixed training data. In OpenCLIP, v1 refers to LAION-COCO NLLB (Visheratin, 2023).
@@
A.6 Prompt Variation
Our evaluation of Large Vision-Language Models (LVLMs) in section˜4 focuses on a single prompt. Given prior work demonstrating the importance of prompt phrasing and advocating for multi-prompt evaluation (Mizrahi et al., 2024; Cheng et al., 2024b), we additionally test a modified version that appears “safer.” This variant appends the instruction “Do not get fooled by typographic attacks.” to the original prompt. Contrary to the expectation formulated in these works, this modification does not reduce vulnerability. As illustrated in fig.˜10, the susceptibility to typographic attacks remains largely unchanged.
@@ A.6 Prompt Variation
A.7 Replacing VLM in LLaVA
Previous work (Cheng et al., 2024a; Li et al., 2024) suggest that susceptibility to typographic attacks can be attributed to a cross-modal alignment deficiency inherent to LVLMs while the functionality of vision encoders is intact. Further, (Malik et al., 2025) demonstrates that integrating an adversarially pretrained vision encoder into LLaVA significantly improves robustness against various attacks, including typographic perturbations introduced in (Li et al., 2024).
Our evaluations on SCAM support the conclusion that some Vision-Language Models (VLMs) exhibit higher typographic robustness than others. In particular, LVLMs using the OpenAI-trained ViT-L-14-336 encoder trained – such as LLaVA (Li et al., 2023b; Liu et al., 2023a, 2024b) – show heightened vulnerability. To isolate the effect of the vision encoder, we train a LLaVA model using the ViT-L-14-CLIPA-336 encoder (X. Li, Z. Wang, and C. Xie (2023c); 13), which was trained on datacomp1b and differs primarily in training data and regime, not architecture. Compared to the OpenAI encoder, the CLIPA model exhibits a substantially lower robustness drop: only 25% from NoSCAM to SCAM, versus 65% (see table˜1). It also outperforms the OpenAI encoder by 3% on average across 38 datasets evaluated using OpenCLIP (Ilharco et al., 2021) [^4] https://github.com/mlfoundations/open_clip/blob/main/docs/openclip_results.csv.
@@ A.7 Replacing VLM in LLaVA
Table˜8 shows that our CLIPA-based LLaVA variant does not outperform the original model on SCAM, despite using a VLM that is more robust in isolation. Performance on other benchmarks evaluated via VLMEvalKit (Duan et al., 2024) also drops. These findings suggest that the benefits of a more robust vision encoder do not transfer under naive LLaVA integration.
@@
Table 8: Performance comparison of the original LLaVA-v1.5-7B and our CLIPA-based LLaVA variant across diverse benchmarks. Despite the CLIPA vision encoder showing stronger robustness in isolation, this does not translate to improved performance on SCAM. Evaluations for the upper section are conducted using VLMEvalKit (Duan et al., 2024). Other results follow the protocol outlined in the main part of the paper.
Benchmark LLaVA-v1.5-7B (orig.) LLaVA-v1.5-7B-CLIPA (ours)
MMBench v1.1 (Liuand et al., 2023) 64.32 60.22
MME Perception (Fu et al., 2023) 1444.81 1378.96
SEED-Bench-1 (img) (Li et al., 2023a) 66.81 64.02
TextVQA (Singh et al., 2019) 46.93 45.57
OCRBench (Liu et al., 2024d) 31.90 31.60
OCR-VQA (Mishra et al., 2019) 60.64 57.39
@@
PAINT (Ilharco et al., 2022) 52.73 53.64
RTA100 (Azuma and Matsui, 2023) 48.30 47.30
A.8 Typographic Robustness vs. Typographic Understanding
The main paper shows that increasing the size of LLM backbones reduces susceptibility to typographic attacks (see results for LLaVA and Gemma3 in table˜1). Simultaneously, benchmarks such as TextVQA and OCRBench, reported on the VLMLeaderboards [^6] https://huggingface.co/spaces/opencompass/open_vlm_leaderboard, indicate that typographic understanding – essentially, reading – also improves with model scale (see table˜9). Together, these findings suggest that typographic robustness emerges as model size increases, alongside improved textual perception. However, (Kimura et al., 2024) reports the opposite correlation between susceptibility and OCR performance in LVLMs, indicating that stronger reading ability alone does not guarantee robustness.
Table 9: Performance of selected LLaVA-Next (Liu et al., 2024b) models on three OCR-related benchmarks (TextVQA (Singh et al., 2019), OCRBench (Liu et al., 2024d) (from VLMEvalKit (Duan et al., 2024)), and OCR-VQA (Mishra et al., 2019)) shows that increasing LLM backbone size leads to improved typographic understanding. All models use the ViT-L-14-336 vision encoder. Results are reported from the OpenVLM leaderboard, evaluated using the VLMEvalKit toolkit (Duan et al., 2024).